# CT-Mamba: A Hybrid Convolutional State Space Model for Low-Dose CT Denoising

### v4

@@ Abstract
Low-dose CT (LDCT) significantly reduces the radiation dose received by patients, however, dose reduction introduces additional noise and artifacts. Currently, denoising methods based on convolutional neural networks (CNNs) face limitations in long-range modeling capabilities, while Transformer-based denoising methods, although capable of powerful long-range modeling, suffer from high computational complexity. Furthermore, the denoised images predicted by deep learning-based techniques inevitably exhibit differences in noise distribution compared to normal-dose CT (NDCT) images, which can also impact the final image quality and diagnostic outcomes. This paper proposes CT-Mamba, a hybrid convolutional State Space Model for LDCT image denoising. The model combines the local feature extraction advantages of CNNs with Mamba’s strength in capturing long-range dependencies, enabling it to capture both local details and global context. Additionally, we introduce an innovative spatially coherent ‘Z’-shaped scanning scheme to ensure spatial continuity between adjacent pixels in the image. We design a Mamba-driven deep noise power spectrum (NPS) loss function to guide model training, ensuring that the noise texture of the denoised LDCT images closely resembles that of NDCT images, thereby enhancing overall image quality and diagnostic value. Experimental results have demonstrated that CT-Mamba performs excellently in reducing noise in LDCT images, enhancing detail preservation, and optimizing noise texture distribution, and exhibits higher statistical similarity with the radiomics features of NDCT images. The proposed CT-Mamba demonstrates outstanding performance in LDCT denoising and holds promise as a representative approach for applying the Mamba framework to LDCT denoising tasks.
@@ 1 Introduction
11footnotetext: Linxuan Li and Wenjia Wei are co-first authors.22footnotetext: Github link: https://github.com/zy2219105/CT-Mamba33footnotetext: Contact Email: linxuan.li@163.com, weiwenjia@buaa.edu.cn
Computed tomography (CT) is an essential imaging technique in clinical practice, providing crucial anatomical information that aids physicians in making appropriate medical decisions [49]. However, frequent CT scans can significantly increase the radiation dose received by patients, particularly in radiation therapy, where multiple scans are often required to achieve precise tumor localization and minimize damage to organs at risk. Excessive radiation may harm the patient, impact organ function, increase the incidence of radiation-related diseases linked to genetic damage, and ultimately reduce the patient’s quality of life. Therefore, reducing the radiation dose of CT imaging has become a focus of attention. Low-dose CT (LDCT) is an effective method for reducing patient radiation exposure [2]. In clinical practice, to acquire LDCT images, it is common to reduce the tube current to decrease the flux of the electron beam that generates X-ray photons, and/or reduce the number of projections during CT procedures. While these methods reduce radiation dose, they also introduce additional noise and artifacts, which decrease the overall quality of the image. If these issues are not effectively addressed, they will significantly impact the application of LDCT in various clinical scenarios. How to obtain high-quality CT image‘s that are comparable to normal-dose CT (NDCT) images and meet the clinical utilities from low-dose scanning protocols has become a long-standing and practically significant problem in the field of CT imaging.
@@
In recent years, with the continuous improvement of computer performance, and the rapid development of cloud computing and distributed computing, deep learning-based image post-processing methods have achieved remarkable achievements in the field of LDCT denoising. Deep learning can automatically learn complex patterns, offering distinct advantages over traditional image post-processing methods. Currently, two popular network architectures, convolutional neural networks (CNNs) [22] and Transformers [37], are widely applied in LDCT denoising. Furthermore, hybrid models combining these two architectures, along with other neural network frameworks, are continually being developed.
CNN-based methods effectively extract image features through convolutional layers, and have achieved remarkable success in tasks such as image restoration [7], object detection [34], and image segmentation [35]. In the field of LDCT denoising, several representative CNN-based models have been proposed [32, 24, 11, 21, 31, 50, 20]. However, CNN frameworks are limited by their lack of long-range modeling capabilities, making it difficult to capture the global contextual information in images [47, 52]. Transformers were originally designed for natural language processing, but have now been successfully adapted for image processing, such as Vision Transformer (ViT) [6] for image recognition and SwinTransformer [27] for a general backbone in various vision tasks. By treating images as sequences of patches, Transformers has better capabilities to capture global information. However, they are computationally expensive, and this is because the self-attention mechanism scales quadratically with the input size [28]. This characteristic creates a serious bottleneck, particularly in high-resolution applications and edge computing scenarios. To overcome these limitations, researchers are working to improve these models.
@@
Currently, Mamba has been rapidly adopted in computer vision tasks. It unfolds image patches into sequences along the horizontal and vertical dimensions of the image, and performs bi-directional scanning along these two directions. This bidirectional scanning method enables VMamba to effectively capture both global and local information within images. At present, some studies have employed the network architectures based on visual Mamba [9, 53, 10, 45, 5]. Additionally, applications in LDCT image denoising have also been reported [16]. The classic visual Mamba scanning methods, such as horizontal and vertical scanning, can cause certain adjacent pixels in the spatial domain to be spaced too far apart when unfolded into a sequence. This separation can undermine Mamba’s performance in visual tasks, particularly in medical imaging, where capturing fine structures and lesions is critical. To address this issue, we have improved the scanning method to ensure the spatial correlation between pixels in the image. Furthermore, by combining Mamba with CNN, we harness CNN’s strengths in local feature extraction along with Mamba’s global modeling capability, allowing the model to excel in handling the complex task of LDCT image denoising.
@@
2. We propose the Coherence Z-Scan State Space Block (CZSS), which introduces an innovative spatially coherent “Z”-shaped scanning scheme. This approach ensures spatial continuity between adjacent pixels in the image, enhancing the model’s ability to preserve details and improve denoising effectiveness.
@@ 2.2 Framework Overview
Figure 3: The structure of Frequency Domain Attention Module (FDAM).
@@ 3.1 Datasets
In this work, we used a publicly released patient dataset for the 2016 NIH-AAPM-Mayo Clinic Low-Dose CT Grand Challenge [30]. This dataset contains paired low-dose and normal-dose abdominal images with a slice thickness of 1 mm for 10 patients. The quarter-dose LDCT images were generated by adding Poisson noise to the projection data of the NDCT images to mimic a noise level that corresponded to 25% of the full dose. We used data from 9 patients for model training, totaling 5,410 image pairs, while data from the remaining patient (L506) was used for testing model performance, totaling 526 pairs. Additionally, using the method in [40], we simulated paired low-dose and normal-dose images for 20 patients from the AMOS dataset (referred to as the Simulator dataset) for model training and validation [18]. The low-dose images also correspond to a quarter of the normal dose. We used data from 19 patients for training, totaling 1,872 image pairs, and data from 1 patient (AMOS-0033) for testing model performance, totaling 106 pairs.
@@ 3.2 Implementation details
The proposed CT-Mamba model is implemented in PyTorch and trained in an environment equipped with an NVIDIA RTX 3090 Ti 24G GPU. The optimizer used is AdamW, with parameters set to ($\beta 1$, $\beta 2$) = (0.9, 0.999) and a weight decay of 0.02. The initial learning rate is set to 1e-3 and is gradually reduced to 1e-6 using cosine annealing. During training, each image is randomly cropped into four 64×64 patches as input for each batch, with the batch size of 8. To balance the numerical relationship between the deep NPS-loss and the L1 loss and ensure training stability, the deep NPS-loss is not introduced during the first 10 epochs. Afterward, the hyperparameters of corresponding weights for the deep NPS loss are set to ${{\lambda}_{2}}$ = (1e-4, 1e-2). The perceptual loss weight ${{\lambda}_{3}}$ is set to 1e-2. The training process lasts for 250 epochs, with the final model selected based on the optimal loss achieved.
@@ 3.3 Experiment result and design
To demonstrate the effectiveness of the proposed CT-Mamba, we selected several advanced representative methods for comparative experiments, including EDCNN [26], RED-CNN [3], Uformer [41], CTformer [38], and VM-UNet [36]. EDCNN and RED-CNN are CNN-based methods, Uformer and CTformer are Transformer-based methods, and VM-UNet is based on the Mamba framework.
@@ 3.3.1 Visual evaluation
This section presents the visual validation results of the CT-Mamba model compared to various baseline methods. In the Mayo dataset, we selected two representative slices, each containing the predicted results from each method along with their difference images relative to the NDCT images, to validate the comprehensive performance of our proposed method in LDCT processing, as shown in Fig. 7 and Fig. 8. To further assess detail representation, an enlarged view of the region of interest (ROI) which is marked by a red dashed rectangle is provided in the lower right corner of each image, along with the PSNR/SSIM quantitative results for each method displayed in the lower left corner.
@@
From the results shown in Fig. 7 and Fig. 8, all methods exhibit a certain degree of noise and artifact removal capability. However, compared to other methods, our proposed approach achieves the optimal processing performance, demonstrating excellent denoising competence in terms of detail preservation and artifact suppression. In the ROIs, we observe that CT-Mamba enhances the representation of fine structures (such as blood vessels), as shown in Fig. 7; low-attenuation lesions are more clearly visible while maintaining well-defined tissue edges, as shown in Fig. 8.
@@
In Fig. 9, we also present the denoising results and difference images of a representative slice from the Simulator dataset, further validating the outstanding denoising performance and generalization capability of CT-Mamba across different datasets. The upper right corner of each predicted image shows the corresponding difference image.
We set the window level of the predicted results to 40 HU and the window width to 400 HU (i.e., a range from -160 HU to 240 HU). For the difference images, the window level is set to 0 HU with a window width of 200 HU (i.e., a range from -100 HU to 100 HU).
Table 1: The average quantitative results of different methods for patient(L506) in the Mayo test dataset.
@@
LDCT 42.3509 0.9471 31.9809 0.2965
EDCNN 46.6959 0.9819 19.1911 0.3578
REDCNN 46.1284 0.9799 20.5049 0.3406
Uformer 44.4342 0.9655 24.8988 0.3192
CTformer 46.2254 0.9770 20.3881 0.3471
VM-Unet 46.5672 0.9824 19.4795 0.3606
CT-Mamba 47.2825 0.9842 17.9850 0.3735
@@
3.3.2 Quantitative analysis
In Table. 1 and Table. 2, we present the average quantitative results for PSNR, SSIM, RMSE, and VIF for patient L506 from the Mayo dataset and patient AMOS-0033 from the Simulator dataset, respectively. A comprehensive evaluation of the PSNR, SSIM, RMSE, and VIF metrics shows that CT-Mamba model outperforms the comparison methods across all quantitative indicators.
@@ 3.3.3 Radiomics reserach
Radiomics holds the potential to transform digital medical images into quantitative features that reveal underlying pathology, with promising applications in tumor classification and patient outcome prediction [44].In this study, the TotalSegmentator [42] extension in 3D Slicer with a pretrained nnU-Net [17] was utilized for automatic segmentation of target organs (aorta, right kidney, liver, stomach, small bowel, and urinary bladder), followed by manual adjustments to ensure accuracy. For patient L506 from the Mayo dataset, 93 first-order radiomics features were extracted for each target organ from the volumes using Pyradiomics (ver. 3.0). An ideal LDCT denoising algorithm is expected to produce LDCT volumes with statistically similar radiomics features to those of NDCT volumes. The following two tests were designed to show statistical similarity:
(1) Statistical Distribution of First-Order Features: In statistical analysis, the Wilcoxon rank-sum test is commonly used to assess whether two datasets share the same distribution. For each volume, the Wilcoxon rank-sum test was employed to evaluate the consistency of radiomics features distribution between the denoised images and the reference NDCT (p $>$ 0.05). In Table 3, we show the p-values for each model across the target organs (aorta, right kidney, liver, stomach, small bowel, and urinary bladder). A higher p-value indicates greater similarity to the radiomics features distribution of the NDCT. As shown in Table 3, while VM-UNet performed well for the liver and stomach, its performance for the aorta was poor, exhibiting significant differences in radiomics features distributios compared to the NDCT. In contrast, the proposed CT-Mamba model demonstrated superior performance across multiple organs, including the aorta, right kidney, and urinary bladder. Furthermore, its radiomics features distribution closely aligned with those of the NDCT for all other target organs. Notably, for the aorta, the CT-Mamba outperformed all other models.
(2) Pairwise Feature Similarity Ratio: To quantitatively assess the similarity between the radiomics features of the target organs for all models and those of the NDCT, pairwise comparisons were performed for the 93 first-order features. The mean absolute error (MAE) between the features of each model and the corresponding NDCT features (10):
@@
where $i$ was the different models, and $V$ was the values of the selected radiomics features. A lower similarity ratio $R_{i}$ indicates higher similarity between the features of the model and those of the NDCT. For each model, the $R$ across the target organs were compared against $R_{\text{LDCT}}$, and the number of features with $R$ exceeding $R_{\text{LDCT}}$ was recorded.
As shown in Table 4, the proposed CT-Mamba demonstrated superior performance across all target organs. Notably, for the small bowel, where other models performed poorly, CT-Mamba achieved the best performance, further highlighting its robustness and effectiveness. The radiomics analysis indicate that the proposed CT-Mamba effectively preserves the shape and texture features of multiple organs within the LDCT images, demonstrating its potential value for clinical applications.
Table 3: The p-values calculated using the Wilcoxon rank-sum test for each model. AP-value of under 0.05 indicates statistical difference. Pairs with statistical differences are marked by *.
Method Aorta Right Kidney Liver Stomach Small Bowel Urinary Bladder
LDCT 0.2774 0.0787 0.1626 0.7841 0.7116 0.6320
EDCNN 0.0557 0.0557 0.3793 0.4701 0.7519 0.5830
REDCNN 0.0979 0.4080 0.7959 0.9191 0.9893 0.6018
Uformer 0.0979 0.1208 0.0386* 0.2137 0.4772 0.1766
CTformer 0.1960 0.1626 0.3544 0.6198 0.6498 0.3040
VM-Unet 0.0016* 0.3519 0.8361 0.9801 0.8286 0.6741
CT-Mamba 0.5014 0.7960 0.6872 0.8948 0.8857 0.8032
Table 4: Number of similarity ratio which are higher for each model compared to the LDCT volumes out of the 93 features.
Method Aorta Right Kidney Liver Stomach Small Bowel Urinary Bladder
EDCNN 6 1 6 32 61 8
REDCNN 5 1 5 33 81 12
Uformer 10 5 5 62 76 10
CTformer 8 1 8 36 63 7
VM-Unet 9 5 6 23 67 13
CT-Mamba 4 2 7 11 4 10

### v5

@@ Abstract
Low-dose CT (LDCT) significantly reduces the radiation dose received by patients, however, dose reduction introduces additional noise and artifacts. Currently, denoising methods based on convolutional neural networks (CNNs) face limitations in long-range modeling capabilities, while Transformer-based denoising methods, although capable of powerful long-range modeling, suffer from high computational complexity. Furthermore, the denoised images predicted by deep learning-based techniques inevitably exhibit differences in noise distribution compared to normal-dose CT (NDCT) images, which can also impact the final image quality and diagnostic outcomes. This paper proposes CT-Mamba, a hybrid convolutional State Space Model for LDCT image denoising. The model combines the local feature extraction advantages of CNNs with Mamba’s strength in capturing long-range dependencies, enabling it to capture both local details and global context. Additionally, we introduce an innovative spatially coherent Z-shaped scanning scheme to ensure spatial continuity between adjacent pixels in the image. We design a Mamba-driven deep noise power spectrum (NPS) loss function to guide model training, ensuring that the noise texture of the denoised LDCT images closely resembles that of NDCT images, thereby enhancing overall image quality and diagnostic value. Experimental results have demonstrated that CT-Mamba performs excellently in reducing noise in LDCT images, enhancing detail preservation, and optimizing noise texture distribution, and exhibits higher statistical similarity with the radiomics features of NDCT images. The proposed CT-Mamba demonstrates outstanding performance in LDCT denoising and holds promise as a representative approach for applying the Mamba framework to LDCT denoising tasks.
@@ 1 Introduction
11footnotetext: Linxuan Li and Wenjia Wei contributed equally to this work.22footnotetext: https://github.com/linxuan-li/CT-Mamba33footnotetext: Email: linxuan.li@163.com, weiwenjia@buaa.edu.cn44footnotetext: DOI: 10.1016/j.compmedimag.2025.102595
Computed tomography (CT) is an essential imaging technique in clinical practice, providing crucial anatomical information that aids physicians in making appropriate medical decisions [58, 32]. However, frequent CT scans can significantly increase the radiation dose received by patients, particularly in radiation therapy, where multiple scans are often required to achieve precise tumor localization and minimize damage to organs at risk. Excessive radiation may harm the patient, impact organ function, increase the incidence of radiation-related diseases linked to genetic damage, and ultimately reduce the patient’s quality of life. Therefore, reducing the radiation dose of CT imaging has become a focus of attention. Low-dose CT (LDCT) is an effective method for reducing patient radiation exposure [2]. In clinical practice, to acquire LDCT images, it is common to reduce the tube current to decrease the flux of the electron beam that generates X-ray photons, and/or reduce the number of projections during CT procedures. While these methods reduce radiation dose, they also introduce additional noise and artifacts, which decrease the overall quality of the image. If these issues are not effectively addressed, they will significantly impact the application of LDCT in various clinical scenarios. How to obtain high-quality CT image‘s that are comparable to normal-dose CT (NDCT) images and meet the clinical utilities from low-dose scanning protocols has become a long-standing and practically significant problem in the field of CT imaging.
@@
In recent years, with the continuous improvement of computer performance, and the rapid development of cloud computing and distributed computing, deep learning-based image post-processing methods have achieved remarkable achievements in the field of LDCT denoising. Compared with traditional post-processing approaches, deep learning models are capable of automatically learning complex patterns in images, thereby exhibiting superior denoising performance. Currently, two popular network architectures, convolutional neural networks (CNNs) [27] and Transformers [44], are widely applied in LDCT denoising. Furthermore, hybrid models combining these two architectures, along with other neural network frameworks, are continually being developed.
CNN-based methods effectively extract image features through convolutional layers, and have achieved remarkable success in tasks such as image denoising [25, 24], object detection [39], and image segmentation [41]. In recent years, CNNs have also been widely applied in medical image processing tasks, such as automated detection of knee joint synovial fluid in MR images [17], multi-class classification of skin lesions in dermoscopic images [18]. In the field of LDCT denoising, several representative CNN-based models have likewise been proposed [37, 30, 10, 26, 36, 59, 23]. However, due to the limited receptive field of convolutional operators, CNNs exhibit certain limitations in modeling long-range dependencies and capturing global semantic information in images. Originally designed for natural language processing, Transformer has been successfully adapted to computer vision tasks, where its self-attention mechanism demonstrates remarkable capabilities in modeling long-range dependencies and capturing global semantic information in images [28]. Vision Transformer was the first to apply this architecture to image classification, achieving performance comparable to CNNs and laying the foundation for the broader adoption of Transformers in visual tasks [6]. Swin Transformer improved computational efficiency by introducing a hierarchical architecture and shifted window attention, gradually evolving into a general-purpose backbone network for various vision tasks. However, Transformer-based models often incur high computational costs due to the quadratic complexity of the self-attention mechanism with respect to the number of input tokens. This characteristic presents a significant bottleneck in high-resolution image processing and edge computing scenarios.
Diffusion models have achieved remarkable results in visual tasks in recent years [55]. Representative works, for example, the Denoising Diffusion Probabilistic Models (DDPM), proposed by Ho et al., define a forward process that gradually corrupts an image by adding Gaussian noise, and train a noise prediction network (typically a U-Net) to approximate the reverse denoising process, thereby progressively reconstructing a clean image from noise [15]. On this basis, various extensions have been proposed to improve modeling efficiency. For instance, Denoising Diffusion Implicit Models (DDIM) accelerate the generation process by constructing a deterministic sampling path with non-Markovian properties [43]. Latent Diffusion Models (LDM) perform the diffusion process in a pretrained low-dimensional latent space, significantly reducing the computational burden for high-resolution image generation [40]. In the field of LDCT denoising, studies have demonstrated the effectiveness of diffusion models in suppressing noise artifacts and improving image quality [52, 33]. However, current diffusion models generally suffer from slow inference speed and high computational cost, which also limit their deployment in scenarios such as edge computing [22, 48]. To overcome these limitations, researchers are working to improve these models.
@@
Currently, Mamba has been rapidly adopted in computer vision tasks. It unfolds image patches into sequences along the horizontal and vertical dimensions of the image, and performs bi-directional scanning along these two directions. This bidirectional scanning method enables VMamba to effectively capture both global and local information within images. At present, some studies have employed the network architectures based on visual Mamba [61, 9, 54, 5]. Additionally, applications in LDCT image denoising have also been reported [16]. However, the classic visual Mamba scanning method (row-column scanning) tends to disrupt the original spatial continuity between adjacent pixels near image boundaries, breaking the spatial adjacency relationships, which leads to local information loss and ultimately impairs the modeling of fine structures [8]. This separation can undermine Mamba’s performance in visual tasks, particularly in medical imaging, where capturing fine structures and lesions is critical. To address this issue, we have improved the scanning method to ensure the spatial correlation between pixels in the image. Furthermore, by combining Mamba with CNN, we harness CNN’s strengths in local feature extraction along with Mamba’s global modeling capability, allowing the model to excel in handling the complex task of LDCT image denoising.
@@
2. We propose the Coherence Z-Scan State Space Block (CZSS), which introduces an innovative spatially coherent Z-shaped scanning scheme. This approach ensures spatial continuity between adjacent pixels in the image, enhancing the model’s ability to preserve details and improve denoising effectiveness.
@@ 2.2 Framework Overview
Figure 3: The structure of Frequency Domain Attention Module (FDAM).
@@ 3.1 Datasets
In this work, we used a publicly released patient dataset for the 2016 NIH-AAPM-Mayo Clinic Low-Dose CT Grand Challenge (referred to as the Mayo dataset) [35]. This dataset contains paired low-dose and normal-dose abdominal images with a slice thickness of 1 mm for 10 patients. The quarter-dose LDCT images were generated by adding Poisson noise to the projection data of the NDCT images to mimic a noise level that corresponded to 25% of the full dose. To ensure the statistical robustness of our quantitative evaluations, we implemented a five-fold cross-validation strategy on the Mayo dataset. Specifically, the 10 patients were divided into 5 groups, with each fold using 2 patients as the test and the remaining 8 for training. For the visual analysis, we used data from 9 patients for model training, totaling 5,410 image pairs, while data from the remaining patient (L506) was used for testing model performance, totaling 526 pairs. Additionally, using the method in [47], we simulated paired low-dose and normal-dose images for 20 patients from the AMOS dataset (referred to as the Simulator dataset) for model training and validation [20]. The low-dose images also correspond to a quarter of the normal dose. We used data from 19 patients for training, totaling 1,872 image pairs, and data from 1 patient (AMOS-0033) for testing model performance, totaling 106 pairs.
We also used in-house real low- and normal-dose head phantom data for experimental validation. Specifically, an anthropomorphic head phantom was scanned using a Cone-Beam Computed Tomography (CBCT) on-board imager (TrueBeam System, Varian Medical Systems, Palo Alto, CA). The phantom data comprises pairs of CT images acquired using low-dose parameters (80 kV, 100 mA) and normal-dose parameters (80 kV, 400 mA), with the low-dose settings representing one-quarter of the radiation exposure of the normal-dose configurations. The data directly reflect the noise and artifact characteristics under real low- and normal-dose conditions in daily CBCT for patient setup in image-guided radiotherapy.
@@ 3.2 Implementation details
The proposed CT-Mamba model is implemented in PyTorch and trained in an environment equipped with an NVIDIA RTX 3090 Ti 24G GPU. The optimizer used is AdamW, with parameters set to ($\beta 1$, $\beta 2$) = (0.9, 0.999) and a weight decay of 0.02. The initial learning rate is set to 1e-3 and is gradually reduced to 1e-6 using cosine annealing. During training, each image is randomly cropped into four 64×64 patches as input for each batch, with the batch size of 8. To balance the numerical relationship between the Deep NPS loss and the L1 loss and ensure training stability, the Deep NPS loss is not introduced during the first 10 epochs. Afterward, the hyperparameters of corresponding weights for the Deep NPS loss are set to ${{\lambda}_{2}}$ = (1e-4, 1e-2). The perceptual loss weight ${{\lambda}_{3}}$ is set to 1e-2. The training process lasts for 250 epochs, with the final model selected based on the optimal loss achieved.
@@ 3.3 Experiment result and design
To demonstrate the effectiveness of the proposed CT-Mamba, we selected several advanced representative methods for comparative experiments, including RED-CNN [3], Uformer [49], CTformer [45], DDPM, and VM-UNet [42]. RED-CNN is CNN-based methods, Uformer and CTformer are Transformer-based methods, DDPM is a diffusion model-based method, and VM-UNet is based on the Mamba framework.
@@ 3.3.1 Visual evaluation
This section presents the visual validation results of the CT-Mamba model compared to various baseline methods. In the Mayo dataset, we selected two representative slices, each containing the predicted results from each method along with their difference images relative to the NDCT images, to validate the comprehensive performance of our proposed method in LDCT processing, as shown in Fig. 7 and Fig. 8. To further assess detail representation, an enlarged view of the ROI which is marked by a red dashed rectangle is provided in the lower right corner of each image, along with the PSNR/SSIM quantitative results for each method displayed in the lower left corner. From the results shown in Fig. 7 and Fig. 8, all methods exhibit a certain degree of noise and artifact removal capability. However, compared to other methods, our proposed approach achieves the optimal processing performance, demonstrating excellent denoising competence in terms of detail preservation and artifact suppression. In the ROIs, we observe that CT-Mamba enhances the representation of fine structures (such as blood vessels), as shown in Fig. 7; low-attenuation lesions are more clearly visible while maintaining well-defined tissue edges, as shown in Fig. 8.
@@
In Fig. 9, we also present the denoised results and difference images of a representative slice from the Simulator dataset, further validating the outstanding denoising performance and generalization capability of CT-Mamba across different datasets. The upper right corner of each predicted image shows the corresponding difference image. Additionally, we selected one representative case from the real dose head phantom data, as shown in Fig. 10. CT-Mamba effectively reduces noise while achieving the best perceptual quality, further demonstrating the generalizability of our method in clinical scenarios.
@@
Figure 10: Visual comparison results on the real low- and normal-dose head phantom data, which comprises pairs of CT images acquired using on-board CBCT scanner with low-dose (80 kV, 100 mA) and normal-dose (80 kV, 400 mA) settings, respectively.
For the Mayo and Simulator datasets, we set the window level of the predicted results to 40 HU and the window width to 400 HU (i.e., a range from -160 HU to 240 HU). For the difference images, the window level is set to 0 HU with a window width of 200 HU (i.e., a range from -100 HU to 100 HU). For the real dose head phantom data, the window level was set to 0 HU and the window width to 800 HU (i.e., a range from -400 HU to 400 HU).
Table 1: Quantitative comparison of various methods on the Mayo dataset under five-fold cross-validation.
@@
LDCT 40.8926±0.8687 0.9309±0.0113 37.8652±3.7472 0.2709±0.0150
RED-CNN 44.8327±1.6669 0.9714±0.0144 24.2221±5.2109 0.3316±0.0231
Uformer 46.1255±0.7708 0.9802±0.0032 20.5797±1.8107 0.3557±0.0167
CTformer 44.5024±0.7586 0.9710±0.0045 24.8519±2.1773 0.3100±0.0144
DDPM 42.8780±1.2888 0.9533±0.0143 29.8810±4.7620 0.3020±0.0151
VM-Unet 45.4553±0.8614 0.9784±0.0035 22.4240±2.4696 0.3417±0.0169
CT-Mamba 46.2251±0.8006 0.9805±0.0033 20.3479±1.8592 0.3587±0.0180
3.3.2 Quantitative analysis
@@
Table 3: The average PSNR, SSIM, RMSE, and VIF results of the real dose head phantom data.
Method PSNR $\uparrow$ SSIM $\uparrow$ RMSE $\downarrow$ VIF $\uparrow$
LDCT 42.7276 0.9703 57.0131 0.2719
RED-CNN 44.0203 0.9809 48.9109 0.2880
Uformer 42.1097 0.8435 60.1177 0.2756
CTformer 44.6989 0.9860 45.1705 0.2862
DDPM 43.7502 0.9408 49.9756 0.2859
VM-Unet 44.0641 0.9857 48.3067 0.2847
CT-Mamba 44.9698 0.9859 43.7079 0.2963
In Table 1 and Table 2, we present the average quantitative results of PSNR, SSIM, RMSE, and VIF obtained from five-fold cross-validation on the Mayo dataset and from patient AMOS-0033 in the Simulator dataset, respectively. A comprehensive evaluation of the PSNR, SSIM, RMSE, and VIF metrics shows that CT-Mamba model outperforms the comparison methods across all quantitative metrics. Similarly, in Table 3, we present the average quantitative results of the real dose head phantom data. The results demonstrate that CT-Mamba maintains excellent performance under real-dose conditions, further demonstrating its potential for clinical application.
@@ 3.3.3 Radiomics reserach
Radiomics holds the potential to transform digital medical images into quantitative features that reveal underlying pathology, with promising applications in tumor classification and patient outcome prediction [53]. In this study, the TotalSegmentator [50] extension in 3D Slicer with a pretrained nnU-Net [19] was utilized for automatic segmentation of target organs (aorta, right kidney, liver, stomach and small bowel), followed by manual adjustments to ensure accuracy. Under the five-fold cross-validation, for all patients in the Mayo dataset, 167 radiomics features (including 18 first-order features, 75 texture features and 74 wavelet-HHH features) were extracted from the volumes of each target organ using PyRadiomics (ver. 3.0). An ideal LDCT denoising algorithm is expected to produce denoised volumes with statistically similar radiomics features to those of NDCT volumes. The following two tests were designed to show statistical similarity:
(1) Statistical Distribution of First-Order Features: In statistical analysis, the Wilcoxon rank-sum test is commonly used to assess whether two datasets share the same distribution. For each volume, the Wilcoxon rank-sum test was employed to evaluate the consistency of radiomics features distribution between the denoised images and the reference NDCT (p $>$ 0.05). In Table 4, we show the p-values for each model across the target organs under five-fold cross-validation. A higher p-value indicates greater similarity to the radiomics features distribution of the NDCT. As shown in Table 4, while CTformer performed well for the liver and small bowel, its performance for the aorta and right kidney was poor, exhibiting significant deviations in radiomics features distributios compared to the NDCT and even underperforming compared to LDCT. In contrast, the proposed CT-Mamba demonstrated superior performance across multiple organs, including the aorta, right kidney, and stomach. Furthermore, its radiomics features distribution closely aligned with those of the NDCT for all other target organs. Notably, for the right kidney, the CT-Mamba outperformed all other models.
Table 4: The p-values calculated using the Wilcoxon rank-sum test for each model under five-fold cross-validation. A p-value of under 0.05 indicates statistical difference. Pairs with statistical differences are marked by *.
Method Aorta Right Kidney Liver Stomach Small Bowel
LDCT 0.3922±0.2562 0.5373±0.3036 0.4044±0.1834 0.4767±0.2833 0.3959±0.2602
RED-CNN 0.4563±0.3196 0.4743±0.3285 0.5146±0.2938 0.5172±0.1997 0.7143±0.2599
Uformer 0.4376±0.3445 0.5783±0.2718 0.6046±0.1714 0.3047±0.2488 0.5771±0.2571
CTformer 0.3579±0.2883 0.4624±0.2734 0.6815±0.1610 0.5447±0.2712 0.7263±0.2022
DDPM 0.0002*±0.0003 0.0005*±0.0005 0.0001*±0.0001 0.0019*±0.0035 0.0009*±0.0014
VM-Unet 0.2805±0.3136 0.4487±0.3539 0.5497±0.2266 0.6527±0.2339 0.6527±0.2339
CT-Mamba 0.4631±0.3179 0.6842±0.2898 0.5722±0.2740 0.6440±0.2634 0.5417±0.2992
(2) Pairwise Feature Similarity Ratio: To quantitatively assess the similarity between the radiomics features of the target organs for all models and those of the NDCT, pairwise comparisons were performed for the all 167 radiomics features. The mean absolute error (MAE) between the features of each model and the corresponding NDCT features (10):
@@
where $i$ was the different models, and $V$ was the values of the selected radiomics features. A lower similarity ratio $R_{i}$ indicates higher similarity between the features of the model and those of the NDCT. For each model, the $R$ across the target organs were compared against $R_{\text{LDCT}}$, and the number of features with $R_{\text{LDCT}}$ exceeding $R$ was recorded.
As shown in Table 5, the proposed CT-Mamba demonstrated superior performance across all target organs. Notably, for the small bowel, where other models exhibited suboptimal performance, CT-Mamba achieved the best performance, further highlighting its robustness and effectiveness. The radiomics analysis indicate that the proposed CT-Mamba effectively preserves the shape and texture features of multiple organs within the LDCT images, demonstrating its potential value for clinical applications.
Table 5: The number of features with lower similarity ratios compared to LDCT (n/167, five-fold cross-validation).
Method Aorta Right Kidney Liver Stomach Small Bowel
RED-CNN 42±26 48±32 45±27 68±30 51±29
Uformer 41±21 45±22 41±24 79±20 69±21
CTformer 50±24 58±26 45±7 67±27 61±12
DDPM 21±12 21±9 17±8 18±8 20±4
VM-Unet 47±23 48±23 45±22 74±13 59±14
CT-Mamba 43±23 43±19 45±26 82±17 75±17